If a car traveling at 28 m/s is brought to a full stop in 4.0 s after the brakes are applied, find the average acceleration during braking


1
Expert's answer
2021-04-13T06:37:31-0400

"v=v_0-at\to a=\frac{v_0-v}{t}=\frac{25-0}{4}=6.25\ (m/s^2)" . Answer